### v2

# Quantize What Counts: Bit Allocation Insights Informed by Spectral Gaps in Keys and Values

Abstract
Large Language Models (LLMs) have introduced significant advancements to the capabilities of Natural Language Processing (NLP) in recent years. However, as these models continue to scale in size, memory constraints pose substantial challenge. Key and Value cache (KV cache) quantization has been well-documented as a promising solution to this limitation. In this work, we provide two novel theorems aimed at enhancing KV quantization methods. Our first theorem, termed Key-Value Norm Disparity, states that the key weight matrices by nature carry richer information compared to the value weight matrices, as evidenced by higher spectral and Frobenius norms across most of the layers. Our second theorem, Key-Driven Quantization, posits that prioritizing the quantization precision of keys over values induces significant improvements to the overall quantization performance. In particular, assigning greater precision to the keys compared to the values achieves a higher degree of precision reduction with minimal impact on model accuracy. We validate these theorems through theory and extensive experiments on several state-of-the-art LLM architectures and benchmarks. These findings offer valuable guidelines for improving KV cache quantization strategies, facilitating more efficient memory utilization without compromising model performance across diverse NLP tasks. Source code is available at https://github.com/mohsenhariri/spectral-kv.
1 Introduction
LLMs have introduced significant advancements to the capabilities of NLP via their ability to express and capture intricate ideas [1, 2], and a key reason for these advancements is model scaling. GPT has seen a tremendous increase across its four versions: GPT-1 at 117M [3], GPT-2 at 1.5B [4], GPT-3 at 175B [5], and GPT-4 at 1.8T parameters [6]. In the open-source space, Llama has grown to 2T [7], Mistral Large is now 123B [8], and DeepSeek V3 is at 671B [9].
However, with this growth in model size comes a prohibitive memory bottleneck in the KV cache [10]. An increase in parameter count requires an increase in context-window length to accommodate chain-of-thought reasoning, which in turn expands the KV cache central to the attention unit. That expansion can impose a restrictive memory bottleneck on the LLM. Already, supported context windows in SOTA systems have grown significantly [11, 12, 13, 14], underscoring the need for a solution. One powerful remedy is quantization, wherein high-precision numbers are reduced to low-precision formats [15, 16]. In LLMs, this can be applied to the KV-cache tensors (e.g., from BF16 to INT4) yielding substantial memory savings with minimal performance loss.
Size K2V4 K4V2
1B 0.06 0.34
8B 0.55 0.75
14B 0.78 0.91
14B 0.76 0.87
Figure 1: Keys carry greater importance. (Left): Keys exhibit higher norms than values across all Llama 3.3-70B layers. (Right): GSM8K accuracy is higher under $\mathrm{K}_{2}\mathrm{V}_{4}$ (2-bit keys, 4-bit values) than $\mathrm{K}_{4}\mathrm{V}_{2}$.
This process is known as KV cache quantization. Extensive work has been done in designing novel KV quantization frameworks, yet none so far have provided deep investigation into mixed-precision schemes, in particular those which focus on the bit allocations split between the Key and Value caches. Instead, previous works have treated the bit allocation as a hyperparameter, seeing it as an object for optimization via grid search rather than taking a deeper dive into the properties of Keys and Values. We provide further motivation for the study in the form of a full description of the current literature, which is detailed in Section 2.
Therefore, there is a present need for a deep exploration into the inherent differences between the Key and the Value caches, as well as a change to the way that quantization is performed. This study aims to fill this need by uncovering fundamental theoretical ideas to improve upon KV cache quantization methods. A preview of these results is illustrated in Figure 1 (left), showing that the spectral norms of the Keys consistently exceed those of the values. Additionally, Figure 1 (right) demonstrates that accuracy across four different models is notably higher for the K4V2 quantization configuration compared to the K2V4 configuration. Our main contributions are summarized as follows:
  • We prove that the expected values of the spectral norm and Frobenius norm are consistently higher for Key weight matrices compared to Value weight matrices (Key-Value Norm Disparity theorem), implying that Key matrices inherently carry more information. We provide detailed empirical validation through observational studies on multiple LLMs.
  • Building on the Key-Value Norm Disparity, we introduce the Key-Driven Quantization theorem, positing that assigning higher bit precision to Keys over Values significantly enhances quantization efficiency. Specifically, this approach allows for greater precision reduction while maintaining minimal accuracy loss.
  • We substantiate the Key-Driven Quantization theorem through two primary forms of evidence: (1) a theoretical analysis explicitly linking it to the previously established Key-Value Norm Disparity, and (2) comprehensive empirical evaluations involving multiple SOTA LLMs, three datasets, and two quantization frameworks.
Section 2 addresses gaps in the existing literature on KV cache quantization, followed by Section 3.1, which introduces the Key-Value Norm Disparity theorem and its supporting evidence. Section 3.2 presents the Key-Driven Quantization theorem along with its corresponding proof, and Section 4 reports the empirical evaluation of both theorems. Finally, Section 5 summarizes the main contributions and discusses their limitations.
2 Related Work
KV Quantization Background.
Quantization frameworks can be categorized in several ways. One way is the time at which the quantization is performed, which is grouped into training-time and post-training (PTQ) quantization techniques [17]. Training-time quantization frameworks can struggle with the usual costs of neural network training, including the dependence on labelled data and longer training times, but these frameworks can achieve more accurate performance due to targeting the weights or activations during the training process. In contrast, post-training quantization techniques are the reverse, struggling with sometimes less accurate performance due to dealing with the values after the training has been completed but achieving lower computation overhead and no longer relying on labelled data which may not be available [16].
There are several settings for PTQ: weight-only, weight-activation, and KV quantization. Weight-only quantization focuses on the weights and achieves strong performance but can struggle with memory overhead [18, 19, 20]. Weight-activation quantization focuses on both the weights and the activations and avoids memory overhead but can struggle with achieving high accuracy [21, 22, 23]. In contrast, KV quantization offers the memory savings seen in weight-activation methods while also maintaining the performance of weight-only quantization [24]. This is due to the KV cache memory overhead inflation caused by longer sequences [25].
KV quantization frameworks can be further categorized into three groups based on the exact quantization scheme being used for the keys and values: outlier redistribution, fixed-precision, and mixed-precision [26]. In outlier redistribution, the outliers in the cache are moved elsewhere or smoothed out. Several outlier redistribution frameworks have been built to varying levels of success [22, 20, 23]. In fixed-precision, the same number of bits is used to quantize certain parts of the cache while in mixed-precision, different numbers of bits are used. Fixed-precision frameworks have also been built to varying levels of success, though they all suffer from a similar problem of neglecting the inherent properties within the KV cache as well as the varying importance of tokens [27, 28, 29].
More focus has been devoted to the third type of KV quantization framework, mixed-precision, but as yet, these works have not looked much into asymmetric allocation into keys and values [30, 24, 31, 32, 33, 34, 35]. KVQuant, for example, focuses on defining an outlier for each vector, such as channel or token, to achieve what the authors term dense-and-sparse quantization. Only KVTuner, QAQ, and SKVQ discuss the asymmetric allocation in some detail, which is described below.
KV Quantization Gaps.
As described above, extensive work has been done in designing new KV quantization frameworks, yet none so far have conducted deeper investigation into mixed-precision schemes. In particular, no studies have focused on making a bit allocation split between the key matrix and the value matrix, nor have they provided a deep theoretical explanation for the inherent differences between the two matrices. Only three studies to date have touched upon this area, with none providing explanations for potential differences between the key and value matrix importance.
The closest study has been a KV quantization framework called KVTuner, which employs layerwise quantization based on the sensitivity of the layers to the KV cache quantization mode and precision pairs [35]. The study makes a brief mention of the difference in attention score errors and perplexity score errors, observing that precision pairs with more bits allocated for keys than for values achieves similar accuracy than for even-split pairs, such as K8V4 versus K8V8. However, the study makes no further analysis into this observation, providing no reason for why this difference is observed. In addition, the quantization approach does not use any underlying insight on this difference between key and values, instead opting for a straightforward grid search over a limited amount of precision pairs such as K8V4, K4V8, K4V2, and so on. Therefore, this allocation split of bits is not the core focus of the framework, with the focus instead being on deploying a layerwise quantization scheme.
Two other studies also touch upon the key-value difference, but like KVTuner, do not provide much depth into the difference. One of these studies is a KV quantization framework called SKVQ, which deploys a sliding window filter along the input tokens to remove tokens that fall outside the window frame and that meet certain criteria [33]. The quantization strategy achieves 2 bit key quantization and 1.5 bit value quantization, but the authors do not explore the possible reasons behind why this difference in bit allocation yielded such powerful results, instead arriving at this bit allocation via hyperparameter search.
The other study is a KV quantization framework called QAQ, which uses the attention output to perform the key quantization and the self-attention output to perform the value quantization [32]. Unlike SKVQ and all other KV quantization strategies, QAQ is the first to target the keys and values in different approaches. However, the difference in the approach lies in the type of data being used to perform the quantization, not in the exact number of bits being allocated to one versus the other. Finally, both SKVQ and QAQ are post-training quantization approaches, making them susceptible to overhead issues during the attention computation.
Figure 2: Frobenius norm for Llama Family. The x-axis represents the layer index in the model while the y-axis represents the Frobenius norm magnitude. The norms are higher for the Key weights than for the Value weights across layers.
3 Dynamics of the Norms of Key and Value
3.1 Key-Value Norm Disparity
Before formalizing our finding, a brief examination is presented on how the magnitudes of the key and value projection matrices influence KV cache quantization. In the self-attention mechanism, the key weights $W^{k}$ scale the token representations that enter the cache, whereas the value weights $W^{v}$ scale the representations that are read out once the attention scores are formed. Because quantization noise is approximately proportional to the dynamic range of the signal, the relative sizes of $\|W^{k}\|$ and $\|W^{v}\|$ determine which part of the cache is more sensitive to a given bit budget. Computing norms across four Llama-3 models (Fig. 2) show that the Frobenius norm of $W^{k}$ consistently exceeds that of $W^{v}$ in almost every layer; the same ordering holds for the spectral norm. Intuitively, this systematic gap motivates the formal statement captured in Theorem Key-Value Norm Disparity below.
Theorem 1 (Key-Value Norm Disparity).
The expectation of the Frobenius norm for the key weight matrices are greater than that for the value weight matrices.
$\mathbb{E}\!\bigl{[}\|W^{K}\|_{F}\bigr{]}\;>\;\mathbb{E}\!\bigl{[}\|W^{V}\|_{F}\bigr{]}.$ (1)
The proof of Theorem 1 is deferred to Appendix B, but its core argument can be sketched as follows: Starting from a standard Xavier initialization[36], the analysis tracks how stochastic-gradient descent updates the Frobenius norms of $W^{k}$ and $W^{v}$. Under isotropic-input assumptions, the cross-term $\langle W^{m},\,\partial\mathcal{L}/\partial W^{m}\rangle_{F}$ becomes negligible, so the expected norm growth is governed by the squared gradient magnitudes. A careful decomposition of those gradients shows that the Key weight update contains an additional factor of the query weight matrix $Q$, whose own norm increases during training. Consequently, the expected squared gradient of $W^{k}$ dominates that of $W^{v}$, causing $\mathbb{E}\bigl{[}\|W^{k}\|_{F}\bigr{]}$ (and, by monotonicity, the spectral norm) to outpace its value counterpart. This asymptotic divergence establishes the persistent norm gap asserted Key-Value Norm Disparity.
Although Fig. 2 illustrates this disparity for the Llama family, supplementary plots for Mistral models in Appendix D confirm that the same pattern holds across diverse architectures [8]. The next subsection investigates how this norm disparity influences quantization error, laying the groundwork for data-driven bit-allocation strategies.
3.2 Key-Driven Quantization
Theorem Key-Value Norm Disparity establishes a persistent norm gap between the Key and Value weight matrices. Since quantization noise scales with the dynamic range of the signal, weight matrices of larger norm inevitably introduce larger absolute errors under identical bit precision. Consequently, assigning equal quantization budgets to the Key and Value caches is sub-optimal.
Consider an additive-residual Transformer block (Layer Norm omitted for clarity [37]),
$h_{l+1}=h_{l}+W_{l}h_{l},$
and let $\Delta_{l}$ denote the quantization perturbation applied to $W_{l}$. A single block incurs
$\|h_{l+1}-\tilde{h}_{l+1}\|_{2}=\|\Delta_{l}h_{l}\|_{2}\leq\|\Delta_{l}\|_{2}\,\|h_{l}\|_{2}.$
After $L$ layers the worst-case deviation satisfies
$\|h_{L}-\tilde{h}_{L}\|_{2}\;\leq\;\Bigl{(}\prod_{i=1}^{L}\|W_{i}\|_{2}\Bigr{)}\;\max_{1\leq i\leq L}\|\Delta_{i}\|_{2}.$
From  Key-Value Norm Disparity:
$\mathbb{E}\!\bigl{[}\|W^{K}\|_{F}^{2}\bigr{]}>\mathbb{E}\!\bigl{[}\|W^{V}\|_{F}^{2}\bigr{]}.$ (i)
Hence any distortion injected on the key path is amplified more strongly throughout the layers.
Let $X\in\mathbb{R}^{\mathrm{seq}\times d_{\mathrm{model}}}$ be the hidden-state matrix at a given layer. The key and value caches are produced by the same input but different projections:
$K=X\,W^{K},\qquad V=X\,W^{V}.$
Because $X$ is common to both paths, multiplying (i) on the left by $X$ and using the sub-multiplicativity of the Frobenius norm, $\|AB\|_{F}\leq\|A\|_{F}\,\|B\|_{2}$, yields
$\mathbb{E}\!\bigl{[}\|K\|_{F}^{2}\bigr{]}>\mathbb{E}\!\bigl{[}\|V\|_{F}^{2}\bigr{]},$ (ii)
i.e. the norm gap persists in the caches themselves.
Quantization error vs. Frobenius norm.
For a matrix $M\in\mathbb{R}^{m\times n}$ the squared Frobenius norm $\|M\|_{F}^{2}$ equals the total energy $\sum_{i,j}M_{ij}^{2}$, or equivalently the sum of eigenvalues of $M^{\top}M$ [^1] Because $\operatorname{tr}(AB)=\operatorname{tr}(BA)$, one has $\operatorname{tr}\!\bigl{(}M^{\top}M\bigr{)}=\operatorname{tr}\!\bigl{(}MM^{\top}\bigr{)}$. The non-zero eigenvalues of $M^{\top}M$ and $MM^{\top}$ are identical and equal to the squared singular values $\sigma_{i}^{2}$ of $M$; hence $\|M\|_{F}^{2}=\sum_{i}\sigma_{i}^{2}=\operatorname{tr}\!\bigl{(}M^{\top}M\bigr{)}$.. When $M$ is quantized to $\tilde{M}$ with $b$ bits, uniform-scalar quantizers yield:
[…]
where [^2] Big-$\Theta$ notation: $f(n)=\Theta(g(n))$ iff $\exists\,c_{1},c_{2}>0,\;n_{0}$ such that $c_{1}g(n)\leq f(n)\leq c_{2}g(n)$ for all $n\geq n_{0}$. the hidden constants depend only on the quantizer variant. Since Key and Value caches share the same shape, minimizing mean square error reduces to minimizing $\|\,\cdot\,\|_{F}^{2}$. When $\|K\|_{F}\gg\|V\|_{F}$ and equal bit-widths are used,
$\mathbb{E}\!\bigl{[}\|K-\tilde{K}\|_{F}^{2}\bigr{]}\;\gg\;\mathbb{E}\!\bigl{[}\|V-\tilde{V}\|_{F}^{2}\bigr{]},$
making the key cache the dominant source of degradation, formally:
Theorem 2 (Key-Driven Quantization).
Let $(b_{K},b_{V})$ denote the bit allocations assigned to the key and value caches, respectively, under a symmetric uniform-scalar quantizer. For any pair $(b_{K},b_{V})$ and $b_{K}>b_{V}$, the expected inference accuracy is strictly higher than for the swapped allocation $(b_{V},b_{K})$, provided that $\mathbb{E}\!\bigl{[}\|K\|_{F}^{2}\bigr{]}>\mathbb{E}\!\bigl{[}\|V\|_{F}^{2}\bigr{]}$.
Figure 3: Singular value distribution of key and value activations for Llama 3.3-70B on C4. The x-axis represents the singular value indices, ordered from the 5th largest to the smallest, while the y-axis denotes the magnitude of the singular values. The shaded region indicates the minimum-to-maximum range across attention heads within each layer, and the solid lines represent the mean singular value magnitude at each index. This zoomed-in view reveals that, beyond the largest singular value (spectral norm), key matrices exhibit consistently higher singular values even in the lower ranges, underscoring their greater representational significance compared to value matrices. For the complete range, see Figure 5 in Appendix E.
Appendix 3 and  4 establishes a norm-dependent upper bound on the quantization error of an arbitrary matrix $M$, and Appendix C.6 applies this bound to show that key caches incur proportionally larger errors than value caches.
To empirically validate the theoretical predictions, we examined the singular value spectra of the K anc V caches across a broad set of models, up to 70B, by extracting every attention head in every layer (Section 4.1). Figure 3 displays the singular values from the fifth largest onward, revealing a systematic separation between keys and values; the complete spectra in Appendix E (Figure 5) show the same trend across all modes. Aggregate norms reinforce this observation: Figures 7 and 7 plot the layer-wise Frobenius and spectral norms of the $K\!V$ cache in Llama-3.3-70B, consistently indicating larger magnitudes for keys than for values.
4 Experimental Results
4.1 Experiment Setup
We conduct two experiments: (i) a quantization-error analysis and (ii) a downstream evaluation. For the quantization-error analysis, we use PyTorch’s weight-packing quantization to measure reconstruction error. For the downstream evaluation, we employ hardware-aligned quantization via Hugging Face’s Optimum Quanto library as back-end (token-wise, 2/4-bit quantization) on Llama-3.2-1B, Llama-3.1-8B [38], Phi-4-14B [39] and DeepSeek-R1-Qwen-14B, and the Half-Quadratic Quantization (HQQ) [40] back-end (channel-wise, 1–8-bit quantization) on Llama-3.1-8B, Llama-3.2-1B, Llama-3.2-3B, and Qwen3-0.6B/1.7B/4B/8B [41]. The quantization-error sweep additionally includes Phi-3-Medium-128K [42], Llama-3.1-Nemotron-70B [43], Mistral-0.3-7B [8], Qwen-2.5-14B and Llama-3.3-70B. See Appendix G for full details on compute resources, software tools, and reproducibility.
Table 1: Quantization errors for KV cache under various configurations. The quantization error is expressed as mean ± std. We use 2-bit, 3-bit, and 4-bit quantization for both $K_{i}$ and $V_{i}$ across three datasets.
Dataset Model 2-bit 3-bit 4-bit
$K_{2}$ $V_{2}$ $K_{3}$ $V_{3}$ $K_{4}$ $V_{4}$
MMLU Llama3.2-1B 4.851 ± 1.037 0.127 ± 0.101 1.037 ± 0.265 0.021 ± 0.015 0.227 ± 0.059 0.005 ± 0.003
Llama3.1-8B-it 6.003 ± 1.782 0.187 ± 0.127 1.082 ± 0.244 0.028 ± 0.019 0.235 ± 0.055 0.006 ± 0.004
Llama3.3-70B-it 4.883 ± 1.106 0.112 ± 0.093 0.942 ± 0.198 0.016 ± 0.012 0.206 ± 0.043 0.003 ± 0.003
Phi4 5.929 ± 1.545 0.657 ± 0.472 1.306 ± 0.231 0.103 ± 0.070 0.286 ± 0.050 0.022 ± 0.015
Mistral0.3-7B 4.718 ± 1.340 0.398 ± 0.405 0.941 ± 0.240 0.059 ± 0.059 0.206 ± 0.053 0.013 ± 0.013
Qwen2.5-14B 5.184 ± 2.241 1.270 ± 1.547 1.005 ± 0.288 0.182 ± 0.221 0.223 ± 0.067 0.040 ± 0.052
DeepSeekR1Q-14B 5.126 ± 2.375 1.406 ± 1.609 0.900 ± 0.269 0.198 ± 0.226 0.199 ± 0.062 0.044 ± 0.052
C4 Llama3.2-1B 4.885 ± 1.056 0.207 ± 0.166 1.074 ± 0.289 0.030 ± 0.024 0.233 ± 0.062 0.006 ± 0.005
Llama3.1-8B-it 6.262 ± 1.789 0.254 ± 0.185 1.128 ± 0.249 0.036 ± 0.026 0.247 ± 0.056 0.008 ± 0.005
Llama3.3-70B-it 4.391 ± 1.027 0.121 ± 0.097 0.847 ± 0.175 0.017 ± 0.013 0.186 ± 0.038 0.004 ± 0.003
Phi4 5.715 ± 1.442 0.850 ± 0.684 1.316 ± 0.245 0.124 ± 0.093 0.291 ± 0.056 0.027 ± 0.020
Mistral0.3-7B 5.027 ± 1.332 0.543 ± 0.493 1.014 ± 0.269 0.079 ± 0.068 0.223 ± 0.060 0.017 ± 0.015
Qwen2.5-14B 4.382 ± 2.170 1.544 ± 1.872 0.846 ± 0.250 0.220 ± 0.265 0.187 ± 0.060 0.048 ± 0.060
DeepSeekR1Q-14B 4.832 ± 2.354 1.651 ± 1.914 0.927 ± 0.283 0.232 ± 0.267 0.201 ± 0.061 0.051 ± 0.060
GSM8K Llama3.2-1B 5.703 ± 1.557 0.179 ± 0.136 1.213 ± 0.352 0.026 ± 0.020 0.266 ± 0.078 0.005 ± 0.004
Llama3.1-8B-it 6.445 ± 1.837 0.213 ± 0.161 1.184 ± 0.268 0.030 ± 0.022 0.257 ± 0.060 0.007 ± 0.005
Llama3.3-70B-it 4.967 ± 1.127 0.113 ± 0.091 0.978 ± 0.203 0.016 ± 0.012 0.214 ± 0.044 0.004 ± 0.003
Phi4 6.610 ± 1.624 0.785 ± 0.598 1.498 ± 0.293 0.116 ± 0.082 0.330 ± 0.064 0.025 ± 0.017
Mistral0.3-7B 5.308 ± 1.367 0.461 ± 0.434 1.065 ± 0.288 0.067 ± 0.061 0.232 ± 0.061 0.015 ± 0.013
Qwen2.5-14B 4.829 ± 2.179 1.736 ± 2.659 0.979 ± 0.264 0.241 ± 0.372 0.214 ± 0.061 0.051 ± 0.077
DeepSeekR1Q-14B 4.477 ± 2.176 1.424 ± 1.752 0.830 ± 0.256 0.200 ± 0.242 0.181 ± 0.058 0.044 ± 0.056
4.2 Quantization Error
In several studies (Table 3 of KIVI[29], Table 8 of GEAR[44], and Table 4 of RotateKV[45]), quantized KV cache representations have been shown to outperform their BF16 full-precision counterparts, an outcome that may seem counterintuitive. Taking BF16 as the gold standard, we attribute this phenomenon to the implicit regularization effect introduced by quantization: by discretizing activation values, quantization injects a small, structured noise into the model’s intermediate representations. Much like dropout or weight decay, this noise can smooth overly sharp decision boundaries and mitigate over-fitting, thereby improving generalization on held-out tasks. To isolate and measure this effect, our first experiment focuses on characterizing the quantization error in the KV cache, without grouping tokens, or using residual tokens in each quantization turn.
In addition to per layer and per head comparison, we average MSE across all heads, layers, and tokens in a batch. This approach yields a global view of how severely each bit-width compresses $\mathbf{K}$ vs. $\mathbf{V}$.
The evaluation uses ten random sequences from C4 [46], MMLU [47], and GSM8K [48], padded to the longest length and generated autoregressively for up to 1 000 tokens. Table 1 presents 2-, 3-, and 4-bit errors for seven representative models (Llama3, Phi4 [39], Mistral, Qwen, DeepSeek-Distilled, and others); the complete results appear in Appendix F. Figure 8 plots the full error curve for Llama-3.3-70B on C4, while Figure 10 shows the per-layer breakdown for Llama-3.1-8B.
4.3 Evaluation
Experiments in Section 4.2 established that K cache incur larger MSE than V cache at every bit-width, suggesting that $\mathbf{K}$ contains proportionally richer information. However, ultimately model quality is measured by downstream metrics such as accuracy rather than reconstruction error. Our second experiments therefore measures how the observed error asymmetry translates into task performance under realistic, deployment-oriented constraints.
KV cache quantization methods rarely de-quantize the entire cache after every generated token because the associated memory traffic defeats the purpose of quantization. Instead, they maintain a residual buffer that holds the most recent tokens in full-precision and flushes it en bloc once full. We replicate this strategy by (i) reserving a fixed-size residual cache of 64 tokens and (ii) grouping activations into 128-element blocks—settings adopted by popular open-source implementations such as KIVI, Flash-Decoding [49] and Marlin [50]. Two hardware-aligned quantization kernels were selected to span the design space:
  • Quanto: token-wise (per-row) quantization applied to the key and value caches, supporting $\{2,4\}$-bit precision.
  • HQQ: channel-wise (per-column) quantization applied to the key and value caches, supporting $\{1,2,4,6,8\}$-bit precision.
For downstream evaluation, we employ three generative datasets that require the model to produce extended outputs rather than select among fixed choices: the GSM8K, the CoQA [51] conversational question answering dataset, and the EQ-Bench [52] long-form structured QA suite.
Results.
Tables 2 and 3 corroborate the norm-based hypothesis by showing that downstream accuracy is far more sensitive to the precision assigned to the key cache than to the value cache.
  • Quanto backend. Across four models and two prompt lengths, allocating $4$ bits to keys and $2$ bits to values ($\mathbf{K}_{4}\mathbf{V}_{2}$) recovers on average $94\%$ of the full-precision score, whereas the inverted scheme ($\mathbf{K}_{2}\mathbf{V}_{4}$) loses up to $30$ percentage points (pp). The margin widens with model capacity: in 1-shot GSM8K, $\mathbf{K}_{4}\mathbf{V}_{2}$ outperforms $\mathbf{K}_{2}\mathbf{V}_{4}$ by $+0.30$ pp on Llama-3.2-1B and by $+0.16$ pp on Phi-4-14B. Notably, $\mathbf{K}_{4}\mathbf{V}_{2}$ attains almost the same accuracy as the symmetric $4$-bit baseline.
  • HQQ backend. The broader sweep over $\{1,2,4,6,8\}$ bits and three datasets paints the same picture. On GSM8K, granting even a single extra bit to keys yields double-digit gains; e.g., Qwen3-8B improves from $0.08$ to $0.54$ accuracy ($+46$ pp) when the extra bit is moved from values to keys. Gaps persist on EQ-Bench and CoQA, where key-centric quantization delivers up to $+62$ pp (Qwen3-0.6B, EQ-Parseable, $1$-bit) and never underperforms the converse setting. Advantages taper but do not vanish at higher precisions: at $6$ bits the key-first policy still adds $7\!-\!12$ pp for 8–14B models.
Table 2: KV cache quantization on GSM8K 1-shot and 8-shot performance employing Optimum Quanto backend. $\mathbf{K}_{i}\mathbf{V}_{j}$ denotes $i$-bit quantization for the K cache and $j$-bit for the V cache using Quanto backend. The $\mathbf{K}_{\mathbf{(4bit)}}\mathbf{V}_{\mathbf{(2bit)}}$ column, highlighted below, shows the optimal bit allocation setting.
1-shot 8-shot
Model $\mathbf{K}_{\mathbf{2}}\mathbf{V}_{\mathbf{2}}$ $\mathbf{K}_{\mathbf{2}}\mathbf{V}_{\mathbf{4}}$ $\mathbf{K}_{\mathbf{4}}\mathbf{V}_{\mathbf{2}}$ $\mathbf{K}_{\mathbf{4}}\mathbf{V}_{\mathbf{4}}$ $\mathbf{K}_{\mathbf{2}}\mathbf{V}_{\mathbf{2}}$ $\mathbf{K}_{\mathbf{2}}\mathbf{V}_{\mathbf{4}}$ $\mathbf{K}_{\mathbf{4}}\mathbf{V}_{\mathbf{2}}$ $\mathbf{K}_{\mathbf{4}}\mathbf{V}_{\mathbf{4}}$
Llama3.2-1B-it 0.033 0.035 0.338 0.357 0.031 0.031 0.289 0.369
Llama3.1-8B-it 0.511 0.547 0.752 0.754 0.408 0.441 0.770 0.782
Phi4-14B 0.759 0.783 0.913 0.923 0.771 0.815 0.927 0.931
DeepSeekR1Q-14B 0.772 0.775 0.865 0.867 0.763 0.792 0.876 0.875
22footnotetext: Although Chain-of-Thought prompting is generally discouraged for reasoning models and can even impair their performance, we deliberately apply 1-shot and 5-shot CoT prompts here to examine the effect of longer input prompts on KV cache quantization.
Table 3: KV cache quantization comparison across GSM8K, EQ-Parseable, and CoQA datasets. The results are obtained on the HQQ quantization backend. Each cell shows accuracies for $\mathbf{K}_{i}\mathbf{V}_{x}$ and $\mathbf{K}_{x}\mathbf{V}_{i}$, where $x$ denotes the mean accuracy across quantization bit-widths $\{1,2,4,6,8\}$, with the percentage-point gap annotated above the comparison sign. See Appendix H for all settings.
GSM8K $\mathbf{K}_{1}\mathbf{V}_{x}$ vs. $\mathbf{K}_{x}\mathbf{V}_{1}$ $\mathbf{K}_{2}\mathbf{V}_{x}$ vs. $\mathbf{K}_{x}\mathbf{V}_{2}$ $\mathbf{K}_{4}\mathbf{V}_{x}$ vs. $\mathbf{K}_{x}\mathbf{V}_{4}$ $\mathbf{K}_{6}\mathbf{V}_{x}$ vs. $\mathbf{K}_{x}\mathbf{V}_{6}$ $\mathbf{K}_{8}\mathbf{V}_{x}$ vs. $\mathbf{K}_{x}\mathbf{V}_{8}$
Qwen3-0.6B $0.00\stackrel{{\scriptstyle\scriptscriptstyle+3\%}}{{<}}\mathbf{0.03}$ $0.00\stackrel{{\scriptstyle\scriptscriptstyle+16\%}}{{<}}\mathbf{0.16}$ $0.04\stackrel{{\scriptstyle\scriptscriptstyle+13\%}}{{<}}\mathbf{0.17}$ $\mathbf{0.34}\stackrel{{\scriptstyle\scriptscriptstyle+16\%}}{{>}}0.18$ $\mathbf{0.34}\stackrel{{\scriptstyle\scriptscriptstyle+16\%}}{{>}}0.18$
Llama-3.2-1B $0.00\stackrel{{\scriptstyle\scriptscriptstyle+5\%}}{{<}}\mathbf{0.05}$ $0.01\stackrel{{\scriptstyle\scriptscriptstyle+18\%}}{{<}}\mathbf{0.19}$ $\mathbf{0.27}\stackrel{{\scriptstyle\scriptscriptstyle+7\%}}{{>}}0.20$ $\mathbf{0.28}\stackrel{{\scriptstyle\scriptscriptstyle+8\%}}{{>}}0.20$ $\mathbf{0.28}\stackrel{{\scriptstyle\scriptscriptstyle+7\%}}{{>}}0.21$
Llama-3.2-3B $0.00\stackrel{{\scriptstyle\scriptscriptstyle+19\%}}{{<}}\mathbf{0.19}$ $0.27\stackrel{{\scriptstyle\scriptscriptstyle+17\%}}{{<}}\mathbf{0.44}$ $\mathbf{0.57}\stackrel{{\scriptstyle\scriptscriptstyle+11\%}}{{>}}0.46$ $\mathbf{0.58}\stackrel{{\scriptstyle\scriptscriptstyle+12\%}}{{>}}0.46$ $\mathbf{0.59}\stackrel{{\scriptstyle\scriptscriptstyle+13\%}}{{>}}0.46$
Qwen3-4B $0.00\stackrel{{\scriptstyle\scriptscriptstyle+41\%}}{{<}}\mathbf{0.41}$ $0.01\stackrel{{\scriptstyle\scriptscriptstyle+49\%}}{{<}}\mathbf{0.50}$ $\mathbf{0.80}\stackrel{{\scriptstyle\scriptscriptstyle+29\%}}{{>}}0.51$ $\mathbf{0.82}\stackrel{{\scriptstyle\scriptscriptstyle+31\%}}{{>}}0.51$ $\mathbf{0.82}\stackrel{{\scriptstyle\scriptscriptstyle+31\%}}{{>}}0.51$
Llama-3.1-8B $0.00\stackrel{{\scriptstyle\scriptscriptstyle+31\%}}{{<}}\mathbf{0.31}$ $0.35\stackrel{{\scriptstyle\scriptscriptstyle+17\%}}{{<}}\mathbf{0.52}$ $\mathbf{0.70}\stackrel{{\scriptstyle\scriptscriptstyle+16\%}}{{>}}0.54$ $\mathbf{0.70}\stackrel{{\scriptstyle\scriptscriptstyle+16\%}}{{>}}0.54$ $\mathbf{0.70}\stackrel{{\scriptstyle\scriptscriptstyle+16\%}}{{>}}0.54$
Qwen3-8B $0.00\stackrel{{\scriptstyle\scriptscriptstyle+48\%}}{{<}}\mathbf{0.48}$ $0.08\stackrel{{\scriptstyle\scriptscriptstyle+46\%}}{{<}}\mathbf{0.54}$ $\mathbf{0.86}\stackrel{{\scriptstyle\scriptscriptstyle+31\%}}{{>}}0.55$ $\mathbf{0.87}\stackrel{{\scriptstyle\scriptscriptstyle+32\%}}{{>}}0.55$ $\mathbf{0.86}\stackrel{{\scriptstyle\scriptscriptstyle+31\%}}{{>}}0.55$
Qwen3-14B $0.00\stackrel{{\scriptstyle\scriptscriptstyle+50\%}}{{<}}\mathbf{0.50}$ $0.26\stackrel{{\scriptstyle\scriptscriptstyle+32\%}}{{<}}\mathbf{0.58}$ $\mathbf{0.87}\stackrel{{\scriptstyle\scriptscriptstyle+27\%}}{{>}}0.60$ $\mathbf{0.87}\stackrel{{\scriptstyle\scriptscriptstyle+27\%}}{{>}}0.60$ $\mathbf{0.87}\stackrel{{\scriptstyle\scriptscriptstyle+27\%}}{{>}}0.60$
Qwen3-32B $0.00\stackrel{{\scriptstyle\scriptscriptstyle+42\%}}{{<}}\mathbf{0.42}$ $0.25\stackrel{{\scriptstyle\scriptscriptstyle+23\%}}{{<}}\mathbf{0.48}$ $\mathbf{0.73}\stackrel{{\scriptstyle\scriptscriptstyle+23\%}}{{>}}0.50$ $\mathbf{0.71}\stackrel{{\scriptstyle\scriptscriptstyle+21\%}}{{>}}0.50$ $\mathbf{0.72}\stackrel{{\scriptstyle\scriptscriptstyle+21\%}}{{>}}0.51$
EQ-Parseable
Qwen3-0.6B $0.00\stackrel{{\scriptstyle\scriptscriptstyle+10\%}}{{<}}\mathbf{10.06}$ $0.00\stackrel{{\scriptstyle\scriptscriptstyle+53\%}}{{<}}\mathbf{52.51}$ $51.23\stackrel{{\scriptstyle\scriptscriptstyle+2\%}}{{<}}\mathbf{52.87}$ $\mathbf{84.21}\stackrel{{\scriptstyle\scriptscriptstyle+32\%}}{{>}}52.63$ $\mathbf{84.91}\stackrel{{\scriptstyle\scriptscriptstyle+33\%}}{{>}}52.28$
Llama-3.2-1B $0.00\stackrel{{\scriptstyle\scriptscriptstyle+34\%}}{{<}}\mathbf{33.68}$ $25.61\stackrel{{\scriptstyle\scriptscriptstyle+37\%}}{{<}}\mathbf{61.52}$ $\mathbf{87.25}\stackrel{{\scriptstyle\scriptscriptstyle+22\%}}{{>}}65.38$ $\mathbf{90.06}\stackrel{{\scriptstyle\scriptscriptstyle+24\%}}{{>}}66.32$ $\mathbf{90.06}\stackrel{{\scriptstyle\scriptscriptstyle+25\%}}{{>}}66.08$
Llama-3.2-3B $0.00\stackrel{{\scriptstyle\scriptscriptstyle+36\%}}{{<}}\mathbf{35.44}$ $68.19\stackrel{{\scriptstyle\scriptscriptstyle+5\%}}{{<}}\mathbf{73.10}$ $\mathbf{90.06}\stackrel{{\scriptstyle\scriptscriptstyle+14\%}}{{>}}76.49$ $\mathbf{89.12}\stackrel{{\scriptstyle\scriptscriptstyle+13\%}}{{>}}76.02$ $\mathbf{89.94}\stackrel{{\scriptstyle\scriptscriptstyle+14\%}}{{>}}76.26$
Qwen3-4B $0.00\stackrel{{\scriptstyle\scriptscriptstyle+54\%}}{{<}}\mathbf{47.60}$ $29.01\stackrel{{\scriptstyle\scriptscriptstyle+33\%}}{{<}}\mathbf{58.01}$ $\mathbf{84.09}\stackrel{{\scriptstyle\scriptscriptstyle+28\%}}{{>}}59.88$ $\mathbf{85.96}\stackrel{{\scriptstyle\scriptscriptstyle+31\%}}{{>}}59.06$ $\mathbf{85.38}\stackrel{{\scriptstyle\scriptscriptstyle+29\%}}{{>}}59.88$
Llama-3.1-8B $0.00\stackrel{{\scriptstyle\scriptscriptstyle+61\%}}{{<}}\mathbf{60.58}$ $\mathbf{77.89}\stackrel{{\scriptstyle\scriptscriptstyle+2\%}}{{>}}75.79$ $\mathbf{95.67}\stackrel{{\scriptstyle\scriptscriptstyle+19\%}}{{>}}77.19$ $\mathbf{96.96}\stackrel{{\scriptstyle\scriptscriptstyle+20\%}}{{>}}77.31$ $\mathbf{97.43}\stackrel{{\scriptstyle\scriptscriptstyle+21\%}}{{>}}77.08$
Qwen3-8B $0.00\stackrel{{\scriptstyle\scriptscriptstyle+62\%}}{{<}}\mathbf{58.48}$ $54.39\stackrel{{\scriptstyle\scriptscriptstyle+17\%}}{{<}}\mathbf{70.41}$ $\mathbf{95.32}\stackrel{{\scriptstyle\scriptscriptstyle+25\%}}{{>}}71.35$ $\mathbf{96.26}\stackrel{{\scriptstyle\scriptscriptstyle+27\%}}{{>}}70.76$ $\mathbf{96.14}\stackrel{{\scriptstyle\scriptscriptstyle+26\%}}{{>}}71.11$
Qwen3-32B $0.00\stackrel{{\scriptstyle\scriptscriptstyle+60\%}}{{<}}\mathbf{47.37}$ $56.14\stackrel{{\scriptstyle\scriptscriptstyle+7\%}}{{<}}\mathbf{61.75}$ $\mathbf{80.12}\stackrel{{\scriptstyle\scriptscriptstyle+22\%}}{{>}}62.46$ $\mathbf{80.47}\stackrel{{\scriptstyle\scriptscriptstyle+23\%}}{{>}}62.34$ $\mathbf{79.77}\stackrel{{\scriptstyle\scriptscriptstyle+22\%}}{{>}}62.57$
CoQA
Qwen3-0.6B $0.20\stackrel{{\scriptstyle\scriptscriptstyle+33\%}}{{<}}\mathbf{0.44}$ $0.32\stackrel{{\scriptstyle\scriptscriptstyle+29\%}}{{<}}\mathbf{0.52}$ $\mathbf{0.65}\stackrel{{\scriptstyle\scriptscriptstyle+16\%}}{{>}}0.53$ $\mathbf{0.69}\stackrel{{\scriptstyle\scriptscriptstyle+23\%}}{{>}}0.53$ $\mathbf{0.69}\stackrel{{\scriptstyle\scriptscriptstyle+23\%}}{{>}}0.53$
Llama-3.2-1B $0.22\stackrel{{\scriptstyle\scriptscriptstyle+29\%}}{{<}}\mathbf{0.42}$ $0.48\stackrel{{\scriptstyle\scriptscriptstyle+13\%}}{{<}}\mathbf{0.57}$ $\mathbf{0.67}\stackrel{{\scriptstyle\scriptscriptstyle+13\%}}{{>}}0.58$ $\mathbf{0.68}\stackrel{{\scriptstyle\scriptscriptstyle+14\%}}{{>}}0.58$ $\mathbf{0.68}\stackrel{{\scriptstyle\scriptscriptstyle+14\%}}{{>}}0.58$
Llama-3.2-3B $0.21\stackrel{{\scriptstyle\scriptscriptstyle+50\%}}{{<}}\mathbf{0.60}$ $\mathbf{0.73}\stackrel{{\scriptstyle\scriptscriptstyle+6\%}}{{>}}0.68$ $\mathbf{0.79}\stackrel{{\scriptstyle\scriptscriptstyle+14\%}}{{>}}0.68$ $\mathbf{0.79}\stackrel{{\scriptstyle\scriptscriptstyle+14\%}}{{>}}0.68$ $\mathbf{0.79}\stackrel{{\scriptstyle\scriptscriptstyle+15\%}}{{>}}0.67$
Qwen3-4B $0.34\stackrel{{\scriptstyle\scriptscriptstyle+34\%}}{{<}}\mathbf{0.62}$ $0.68\stackrel{{\scriptstyle\scriptscriptstyle+3\%}}{{<}}\mathbf{0.70}$ $\mathbf{0.81}\stackrel{{\scriptstyle\scriptscriptstyle+12\%}}{{>}}0.71$ $\mathbf{0.81}\stackrel{{\scriptstyle\scriptscriptstyle+12\%}}{{>}}0.71$ $\mathbf{0.81}\stackrel{{\scriptstyle\scriptscriptstyle+13\%}}{{>}}0.70$
Llama-3.1-8B $0.21\stackrel{{\scriptstyle\scriptscriptstyle+50\%}}{{<}}\mathbf{0.60}$ $\mathbf{0.72}\stackrel{{\scriptstyle\scriptscriptstyle+7\%}}{{>}}0.66$ $\mathbf{0.78}\stackrel{{\scriptstyle\scriptscriptstyle+14\%}}{{>}}0.67$ $\mathbf{0.78}\stackrel{{\scriptstyle\scriptscriptstyle+15\%}}{{>}}0.67$ $\mathbf{0.78}\stackrel{{\scriptstyle\scriptscriptstyle+14\%}}{{>}}0.67$
Qwen3-8B $0.38\stackrel{{\scriptstyle\scriptscriptstyle+37\%}}{{<}}\mathbf{0.68}$ $\mathbf{0.73}\stackrel{{\scriptstyle\scriptscriptstyle+1\%}}{{>}}0.72$ $\mathbf{0.82}\stackrel{{\scriptstyle\scriptscriptstyle+12\%}}{{>}}0.72$ $\mathbf{0.82}\stackrel{{\scriptstyle\scriptscriptstyle+12\%}}{{>}}0.72$ $\mathbf{0.82}\stackrel{{\scriptstyle\scriptscriptstyle+12\%}}{{>}}0.72$
Qwen3-14B $0.32\stackrel{{\scriptstyle\scriptscriptstyle+45\%}}{{<}}\mathbf{0.69}$ $\mathbf{0.79}\stackrel{{\scriptstyle\scriptscriptstyle+8\%}}{{>}}0.72$ $\mathbf{0.82}\stackrel{{\scriptstyle\scriptscriptstyle+12\%}}{{>}}0.72$ $\mathbf{0.82}\stackrel{{\scriptstyle\scriptscriptstyle+12\%}}{{>}}0.72$ $\mathbf{0.82}\stackrel{{\scriptstyle\scriptscriptstyle+12\%}}{{>}}0.72$
Qwen3-32B $0.32\stackrel{{\scriptstyle\scriptscriptstyle+44\%}}{{<}}\mathbf{0.68}$ $\mathbf{0.79}\stackrel{{\scriptstyle\scriptscriptstyle+9\%}}{{>}}0.72$ $\mathbf{0.82}\stackrel{{\scriptstyle\scriptscriptstyle+12\%}}{{>}}0.73$ $\mathbf{0.82}\stackrel{{\scriptstyle\scriptscriptstyle+12\%}}{{>}}0.73$ $\mathbf{0.82}\stackrel{{\scriptstyle\scriptscriptstyle+12\%}}{{>}}0.73$
5 Conclusion
Large scale language models increasingly spend the majority of their inference footprint on the KV cache, making cache compression a decisive factor in practical deployment. This work supplied a principled answer to the bit-allocation problem by unifying theory and experiment.
  • Key–Value Norm Disparity. We proved that Key weights exhibit larger Frobenius norms than Value weights throughout training, implying a higher information density in the key path.
  • Key-Driven Quantization. Leveraging this norm gap, a second theorem shows that assigning the finer bit-width to keys and the coarser bit-width to values strictly minimises quantization error, thereby preserving downstream accuracy.
  • Empirical Validation. Experiments across nine model families, three generative benchmarks, and two hardware-aligned back-ends confirm the theory: a $4\!\times\!2$ (keys / values) split consistently recovers $\geq 94\%$ of full-precision accuracy while halving KV-cache memory.
These findings elevate bit allocation from a hyper-parameter tuned by grid search to a design principle derived from model geometry, offering immediate guidance for practitioners and a foundation for hardware–algorithm co-design. We used a context length of 2K tokens, a common setting, but will use longer contexts in future.
Appendix
Appendix A Dynamics of the Norms of Key and Value Weight Matrices
A.1 Preliminary: Dimensions
$\begin{array}[]{@{}ll@{}}c&\text{sequence length (context size)},\\[4.0pt] d_{m}&\text{embedding dimension},\\[4.0pt] d_{v}=d_{k}=d_{m}/n_{h}=d_{m}&\text{for a single head ($n_{h}=1$)},\\[4.0pt] n_{h}&\text{number of attention heads}.\end{array}$
A.2 Symbols
$\begin{array}[]{@{}ll@{}}X&\text{input matrix }(c\times d_{m}),\\ X^{O}&\text{output matrix }(c\times d_{m}),\\ W^{V}&\text{value weights }(d_{m}\times d_{m}),\\ W^{K}&\text{key weights }(d_{m}\times d_{m}),\\ W^{Q}&\text{query weights }(d_{m}\times d_{m}),\\ K=XW^{K}&(c\times d_{m}),\\ Q=XW^{Q}&(c\times d_{m}),\\ V=XW^{V}&(c\times d_{m}),\\ S=QK^{\top}/\sqrt{d_{m}}&(c\times c),\\ A=\operatorname{softmax}(S)&(c\times c),\\ \mathcal{L}&\text{loss function.}\end{array}$
Appendix B Norm Dynamics
In this appendix, we compare the long–time behavior of the Frobenius norms of the key ($W^{K}$) and value ($W^{V}$) weight matrices of a single–head self–attention layer trained with stochastic gradient descent (SGD). We show that, under standard isotropic assumptions, $\|W^{K}\|_{F}$ grows faster than $\|W^{V}\|_{F}$.
B.1 Update rule
At update step $i$ the weights follow
$W^{m}_{i+1}\;=\;W^{m}_{i}-\eta\,\frac{\partial\mathcal{L}}{\partial W^{m}_{i}},\qquad m\in\{K,V\},$ (2)
with learning rate $\eta$. Squaring the Frobenius norm of both sides gives
$\|W^{m}_{i+1}\|_{F}^{2}$ $=\|W^{m}_{i}\|_{F}^{2}+\eta^{2}\Bigl{\|}\frac{\partial\mathcal{L}}{\partial W^{m}_{i}}\Bigr{\|}_{F}^{2}-2\eta\,\Bigl{\langle}W^{m}_{i},\frac{\partial\mathcal{L}}{\partial W^{m}_{i}}\Bigr{\rangle}_{F},$ (3)
where
$\|A\|_{F}$ $\equiv\sqrt{\operatorname{tr}(A^{\top}A)},$ (4)
$\langle A,B\rangle_{F}$ $\equiv\operatorname{tr}(A^{\top}B).$ (5)
B.2 Expectation over mini-batches
Taking an expectation over mini-batches yields
$\Delta\mathbb{E}\bigl{[}\|W^{m}_{i}\|_{F}^{2}\bigr{]}$ $\equiv\mathbb{E}\bigl{[}\|W^{m}_{i+1}\|_{F}^{2}\bigr{]}-\mathbb{E}\bigl{[}\|W^{m}_{i}\|_{F}^{2}\bigr{]}$ (6)
[…]
In high–dimensional weight space the gradient is almost orthogonal to the current weights, making the second term negligible. Hence,
$\Delta\mathbb{E}\bigl{[}\|W^{m}_{i}\|_{F}^{2}\bigr{]}\;\approx\;\eta^{2}\,\mathbb{E}\Bigl{[}\Bigl{\|}\frac{\partial\mathcal{L}}{\partial W^{m}_{i}}\Bigr{\|}_{F}^{2}\Bigr{]}.$ (8)
B.3 Initialization
Throughout we assume Xavier initialization: each entry of $W^{K}$ and $W^{V}$ is drawn i.i.d. from a zero–mean distribution whose variance preserves the input scale [36].
B.4 Gradients
The attention outputs are
$X^{O}\;=\;A\,V\,W^{O},$ (9)
where $A=\operatorname{softmax}\!\bigl{(}QK^{\top}/\sqrt{d_{m}}\bigr{)}$ and $V=XW^{V}$. Differentials give
$d\mathcal{L}(X^{O})$ $=\operatorname{tr}\!\bigl{[}(\partial\mathcal{L}/\partial X^{O})^{\!\top}\,AX\,dW^{V}W^{O}\bigr{]}=\operatorname{tr}\!\bigl{[}(AX)^{\top}\,\frac{\partial\mathcal{L}}{\partial X^{O}}\,(W^{O})^{\top}\,dW^{V}\bigr{]},$ (10)
so that
$\frac{\partial\mathcal{L}}{\partial W^{V}}=(AX)^{\top}\,\frac{\partial\mathcal{L}}{\partial X^{O}}\,(W^{O})^{\top}.$ (11)
Similarly, writing $S=QK^{\top}/\sqrt{d_{m}}$,
$d\mathcal{L}(S)$ $=\operatorname{tr}\!\bigl{[}(\partial\mathcal{L}/\partial S)^{\!\top}\,dS\bigr{]}=\frac{1}{\sqrt{d_{m}}}\,\operatorname{tr}\!\bigl{[}X^{\top}\,(\partial\mathcal{L}/\partial S)^{\!\top}\,Q\,dW^{K}\bigr{]},$ (12)
yielding
$\frac{\partial\mathcal{L}}{\partial W^{K}}=\frac{1}{\sqrt{d_{m}}}\,X^{\top}\,(\partial\mathcal{L}/\partial S)^{\!\top}\,Q.$ (13)
B.5 Expected gradient norms
Squaring the Frobenius norm of (11) and taking an expectation under the isotropic-input assumption,
$\mathbb{E}\!\bigl{[}\|\partial\mathcal{L}/\partial W^{V}\|^{2}_{F}\bigr{]}$ $=c\,\sigma_{x}^{2}\,\sigma_{o}^{2},$ (14)
where $\mathbb{E}[X^{\top}X]=c\sigma_{x}^{2}I$ and $\mathbb{E}\Bigl{[}\Bigl{(}A^{\top}\tfrac{\partial\mathcal{L}}{\partial X^{O}}(W^{O})^{\top}\Bigr{)}\Bigl{(}A^{\top}\tfrac{\partial\mathcal{L}}{\partial X^{O}}(W^{O})^{\top}\Bigr{)}^{\top}\Bigr{]}=\sigma_{o}^{2}I$.
For the key weights,
$\mathbb{E}\!\bigl{[}\|\partial\mathcal{L}/\partial W^{K}\|^{2}_{F}\bigr{]}=\frac{\sigma_{x}^{2}\,\sigma_{s}^{2}}{d_{m}}\,\|Q\|_{F}^{2},$ (15)
with $\sigma_{s}^{2}$ the entry-wise variance of $\partial\mathcal{L}/\partial S$.
B.6 Asymptotic comparison
Substituting these expectations into (8) shows that
$\mathbb{E}\!\bigl{[}\|W^{K}\|^{2}_{F}\bigr{]}\text{grows as}\;\eta^{2}\,\frac{\sigma_{x}^{2}\sigma_{s}^{2}}{d_{m}}\,\|Q\|_{F}^{2},\quad\text{whereas}\quad\mathbb{E}\!\bigl{[}\|W^{V}\|^{2}_{F}\bigr{]}\;\text{grows as}\;\eta^{2}\,c\sigma_{x}^{2}\sigma_{o}^{2}.$ (16)
Because $\|Q\|_{F}^{2}$ itself increases during training, $\|W^{K}\|_{F}$ eventually dominates:
$\mathbb{E}\!\bigl{[}\|W^{K}\|_{F}\bigr{]}\;>\;\mathbb{E}\!\bigl{[}\|W^{V}\|_{F}\bigr{]}.$ (17)
As shown in Equation 17, the expectation of the Frobenius norm of $W^{k}$ is greater than that of $W^{v}$. [^3] While this analysis assumes a single-head, decoder-only transformer for clarity, the extension to multi-head attention is direct. The same conclusion also holds for group-query and multi-query attention.
Appendix C Quantization and Dequantization Error Bounds
In this appendix we establishes upper bounds on the quantization error incurred when a matrix is represented using a finite bit-width in two’s-complement format. Two theorems are presented: one that characterizes the error in terms of the spectral norm and another in terms of the Frobenius norm.
The first theorem demonstrates that the spectral norm of the quantization error is bounded by
$\|A-\widehat{A}\|_{2}\;\lesssim\;\frac{\sqrt{m\,n}}{2^{b}}\,\|A\|_{2}.$
This result implies that, for a given bit-width $b$, matrices with larger spectral norms incur proportionally larger quantization errors. Consequently, a matrix that exhibits a larger spectral norm is more susceptible to quantization errors. To control the error propagation in such matrices, a higher bit-width is necessary.
The second theorem provides an analogous bound for the Frobenius norm,
$\|A-\widehat{A}\|_{F}\;\lesssim\;\frac{\sqrt{m\,n}}{2^{b}}\,\|A\|_{F}.$
Similar to the spectral norm result, this bound indicates that the quantization error, measured in the Frobenius norm, is directly proportional to the norm of the original matrix. Hence, matrices with larger Frobenius norms are also more vulnerable to quantization errors and would benefit from a higher precision during quantization.
C.1 Preliminaries and Notation
Let $A\in\mathbb{R}^{m\times n}$ denote a real matrix whose entries are to be stored using a fixed number of bits in two’s-complement representation. The following notation is adopted:
  • Bit Depth. Given $b$ bits in two’s-complement format, each representable integer $q$ lies in the interval
    $q\in\{-2^{b-1},-2^{b-1}+1,\dots,2^{b-1}-1\}.$
  • Maximum Entry Magnitude. Define
[…]
  • Scale Factor. Set
    $\alpha\;=\;\frac{M}{2^{b-1}-1}.$
    This choice ensures that the scaled entries $A_{ij}/\alpha$ lie within the representable range.
  • Quantization. Define the integer matrix $Q\in\mathbb{Z}^{m\times n}$ by
    $Q_{ij}$ $=\mathrm{round}\!\Bigl{(}\frac{A_{ij}}{\alpha}\Bigr{)},$ (18)
    $\text{with}\quad Q_{ij}$$\in\{-2^{b-1},\dots,2^{b-1}-1\}.$
  • Dequantization (Reconstruction). The reconstructed matrix $\widehat{A}$ is given by
    $\widehat{A}_{ij}\;=\;\alpha\,Q_{ij}.$
Objective. The aim is to bound the errors
$\|A-\widehat{A}\|_{2}\quad\text{and}\quad\|A-\widehat{A}\|_{F},$
in terms of $b$, $m$, $n$, and the norms of $A$.
C.2 Spectral Norm Bound
Theorem 3 (Spectral Norm Quantization Error).
Let $A\in\mathbb{R}^{m\times n}$ and $b\in\mathbb{N}$ be given. Consider two’s-complement quantization with scale factor
$\alpha\;=\;\frac{M}{2^{b-1}-1},\quad M\;=\;\max_{i,j}|A_{ij}|.$
Define
$Q_{ij}\;=\;\mathrm{round}\!\Bigl{(}\frac{A_{ij}}{\alpha}\Bigr{)}\quad\text{and}\quad\widehat{A}_{ij}\;=\;\alpha\,Q_{ij}.$
Then, the following bound holds:
$M$ $\leq\|A-\widehat{A}\|_{2}\leq\sqrt{m\,n}\,\frac{M}{2(2^{b-1}-1)}$ (19)
$\leq\sqrt{m\,n}\,\frac{\|A\|_{2}}{2(2^{b-1}-1)}.$
In approximate form for large $b$,
$\|A-\widehat{A}\|_{2}\;\lesssim\;\frac{\sqrt{m\,n}}{2^{b}}\,\|A\|_{2}.$
Proof.
Entrywise Bound. By construction,
$\left|\frac{A_{ij}}{\alpha}-Q_{ij}\right|\;\leq\;\frac{1}{2}.$
Multiplying by $\alpha$ gives
$|A_{ij}-\widehat{A}_{ij}|\;\leq\;\frac{\alpha}{2}\;=\;\frac{M}{2\,(2^{b-1}-1)}.$
Thus,
$\max_{i,j}\,|A_{ij}-\widehat{A}_{ij}|\;\leq\;\frac{M}{2\,(2^{b-1}-1)}.$
Conversion to the Spectral Norm. Using the inequality
$\|B\|_{2}\;\leq\;\sqrt{m\,n}\,\max_{i,j}|B_{ij}|,$
with $B=A-\widehat{A}$, it follows that
$\|A-\widehat{A}\|_{2}\;\leq\;\sqrt{m\,n}\,\frac{M}{2\,(2^{b-1}-1)}.$
Relating $M$ to $\|A\|_{2}$. Since
$M\;=\;\max_{i,j}|A_{ij}|\;\leq\;\|A\|_{2},$
the bound can be written as
$\|A-\widehat{A}\|_{2}\;\leq\;\sqrt{m\,n}\,\frac{\|A\|_{2}}{2\,(2^{b-1}-1)}.$
For large $b$, where $2^{b-1}-1\approx 2^{b-1}$, the bound becomes
$\|A-\widehat{A}\|_{2}\;\lesssim\;\frac{\sqrt{m\,n}}{2^{b}}\,\|A\|_{2}.$
C.3 Frobenius Norm Bound
Theorem 4 (Frobenius Norm Quantization Error).
Under the same setup, the Frobenius norm of the quantization error satisfies
$\begin{split}\|A-\widehat{A}\|_{F}&\;\leq\;\sqrt{m\,n}\,{\textstyle\frac{M}{2(2^{b-1}-1)}}\\ &\;\leq\;\sqrt{m\,n}\,{\textstyle\frac{\|A\|_{F}}{2(2^{b-1}-1)}}.\end{split}$ (20)
In approximate form,
$\|A-\widehat{A}\|_{F}\;\lesssim\;\frac{\sqrt{m\,n}}{2^{b}}\,\|A\|_{F}.$
Proof.
Entrywise Bound. As established,
$|A_{ij}-\widehat{A}_{ij}|\;\leq\;\frac{M}{2\,(2^{b-1}-1)}\quad\text{for all }i,j.$
Conversion to the Frobenius Norm. By definition,
$\|A-\widehat{A}\|_{F}^{2}\;=\;\sum_{i=1}^{m}\sum_{j=1}^{n}(A_{ij}-\widehat{A}_{ij})^{2},$
which yields
$\|A-\widehat{A}\|_{F}^{2}\;\leq\;m\,n\left(\frac{M}{2\,(2^{b-1}-1)}\right)^{2}.$
Taking square roots leads to
$\|A-\widehat{A}\|_{F}\;\leq\;\sqrt{m\,n}\,\frac{M}{2\,(2^{b-1}-1)}.$
Relating $M$ to $\|A\|_{F}$. Since
$M^{2}\;\leq\;\sum_{i=1}^{m}\sum_{j=1}^{n}A_{ij}^{2}\;=\;\|A\|_{F}^{2},$
it follows that $M\leq\|A\|_{F}$ and hence
$\|A-\widehat{A}\|_{F}\;\leq\;\sqrt{m\,n}\,\frac{\|A\|_{F}}{2\,(2^{b-1}-1)}.$
For large $b$, this simplifies to
$\|A-\widehat{A}\|_{F}\;\lesssim\;\frac{\sqrt{m\,n}}{2^{b}}\,\|A\|_{F}.$
Remark. The results indicate that both the spectral norm and Frobenius norm errors satisfy similar approximate bounds:
$\|A-\widehat{A}\|_{2}$ $\lesssim\frac{\sqrt{m\,n}}{2^{b}}\,\|A\|_{2},$ (21)
$\|A-\widehat{A}\|_{F}$ $\lesssim\frac{\sqrt{m\,n}}{2^{b}}\,\|A\|_{F}.$
C.4 Quantizing KV Cache
Consider the key and value cache
$\mathbf{V}\in\mathbb{R}^{L\times d_{\text{head}}},\quad\mathbf{K}\in\mathbb{R}^{L\times d_{\text{head}}},$
with quantization bit-widths denoted by $b_{V}$ and $b_{K}$, respectively. Let $\widehat{\mathbf{V}}$ and $\widehat{\mathbf{K}}$ denote the dequantized matrices, and define the quantization errors as
$\mathbf{E}_{V}\;=\;\mathbf{V}-\widehat{\mathbf{V}},\quad\mathbf{E}_{K}\;=\;\mathbf{K}-\widehat{\mathbf{K}}.$
An empirical observation is that
$\|\mathbf{V}\|_{\ast}\;<\;\|\mathbf{K}\|_{\ast},$
where $\ast$ denotes either the spectral norm ($\|\cdot\|_{2}$) or the Frobenius norm ($\|\cdot\|_{F}$). In practice, $\mathbf{K}$ typically exhibits a larger norm than $\mathbf{V}$.
C.5 Spectral-Norm Perspective
Standard quantization error bounds yield
$\|\mathbf{E}_{V}\|_{2}$ $\lesssim\frac{\sqrt{L\,d_{\text{head}}}}{2^{b_{V}}}\;\|\mathbf{V}\|_{2},$ (22)
$\|\mathbf{E}_{K}\|_{2}$ $\lesssim\frac{\sqrt{L\,d_{\text{head}}}}{2^{b_{K}}}\;\|\mathbf{K}\|_{2}.$
To achieve comparable spectral-norm errors (i.e., $\|\mathbf{E}_{V}\|_{2}\approx\|\mathbf{E}_{K}\|_{2}$), it is necessary that
$\frac{\sqrt{L\,d_{\text{head}}}}{2^{b_{V}}}\;\|\mathbf{V}\|_{2}\;\approx\;\frac{\sqrt{L\,d_{\text{head}}}}{2^{b_{K}}}\;\|\mathbf{K}\|_{2}.$
Cancelling the common factor $\sqrt{L\,d_{\text{head}}}$ yields
[…]
or equivalently,
$2^{\,b_{K}-b_{V}}\;\approx\;\frac{\|\mathbf{V}\|_{2}}{\|\mathbf{K}\|_{2}}.$
Since $\|\mathbf{V}\|_{2}<\|\mathbf{K}\|_{2}$, it follows that $b_{K}>b_{V}$.
C.6 Frobenius-Norm (MSE) Perspective
Similarly, the Frobenius norm error bounds are given by
$\|\mathbf{E}_{V}\|_{F}$ $\;\lesssim\;\frac{\sqrt{L\,d_{\text{head}}}}{2^{b_{V}}}\;\|\mathbf{V}\|_{F},$ (23)
$\|\mathbf{E}_{K}\|_{F}$ $\;\lesssim\;\frac{\sqrt{L\,d_{\text{head}}}}{2^{b_{K}}}\;\|\mathbf{K}\|_{F}.$
To achieve comparable Frobenius errors,
$\frac{\sqrt{L\,d_{\text{head}}}}{2^{b_{V}}}\;\|\mathbf{V}\|_{F}\;\approx\;\frac{\sqrt{L\,d_{\text{head}}}}{2^{b_{K}}}\;\|\mathbf{K}\|_{F},$
which implies
[…]
and therefore,
$2^{\,b_{K}-b_{V}}\;\approx\;\frac{\|\mathbf{V}\|_{F}}{\|\mathbf{K}\|_{F}}.$
Again, because $\|\mathbf{V}\|_{F}<\|\mathbf{K}\|_{F}$, it follows that $b_{K}>b_{V}$.
Appendix D Supplementary Key-Value Weight Norm Analysis
Here we present additional evidence supporting the Key-Value Norm Disparity theorem established in Section 3.1. While Fig. 2 demonstrates this observation for the Llama model family, the same pattern consistently emerges across Mistral family.
Figure 4: Frobenius norm plot for Mistral Family. The x-axis represents the layer index in the model while the y-axis represents the Frobenius norm magnitude. The spectral norms are higher for the key weights than for the value weights across layers.
Appendix E Extended Visualization of Singular Value Distributions
Figure 5 provides the full-range view of the singular value distribution for the key and value cache at various layers of the Llama 3.3 70B model on the C4 dataset. The horizontal axis enumerates the singular values in descending order, beginning with the largest (i.e., the spectral norm), and the vertical axis measures the corresponding magnitudes. The shaded region captures the minimum-to-maximum range of singular values across the attention heads within each layer, while the solid curves depict the mean singular value at each rank.
Figure 5: Complete singular value distribution of key and value cache for the Llama 3.3 70B model on the C4 dataset. The x-axis denotes the singular value indices, ordered from the largest (spectral norm) to the smallest, while the y-axis represents the corresponding magnitudes. The shaded region illustrates the range between the minimum and maximum singular values across attention heads within each layer, and the solid lines indicate the mean singular value magnitude at each index. This full-spectrum view highlights that key matrices consistently maintain significantly higher singular values throughout the entire distribution, further reinforcing their dominant representational capacity compared to value matrices.
Figure 6: Frobenius norm of key and value cache across layers for the Llama 3.3 70B model on the C4 dataset. The x-axis represents the layer index. The shaded region shows the min-max range across attention heads within each layer, and the solid line indicates the mean norm per layer. Key matrices exhibit significantly higher norms compared to value matrices, highlighting their stronger representational capacity.
Figure 7: Spectral norm of key and value cache across layers for the Llama 3.3 70B model on the C4 dataset. The x-axis represents the layer index. The shaded region shows the min-max range across attention heads within each layer, and the solid line indicates the mean norm per layer. Key matrices exhibit significantly higher norms compared to value matrices, highlighting their stronger representational capacity.
Appendix F Extended Quantization Error Results
Table 4: Quantization error (mean ± std) for the key and value caches ($K_{i}$ and $V_{i}$) at 2-bit, 3-bit, and 4-bit quantization, evaluated on the MMLU dataset.
Model K2 V2 K3 V3 K4 V4
Llama3.2-1B 4.851 ± 1.037 0.127 ± 0.101 1.037 ± 0.265 0.021 ± 0.015 0.227 ± 0.059 0.005 ± 0.003
Llama3.2-1B-it 4.373 ± 1.034 0.124 ± 0.090 0.879 ± 0.218 0.019 ± 0.013 0.192 ± 0.047 0.004 ± 0.003
Llama3.2-3B 3.943 ± 0.924 0.193 ± 0.096 0.849 ± 0.150 0.030 ± 0.015 0.183 ± 0.031 0.007 ± 0.003
Llama3.2-3B-it 4.487 ± 1.180 0.202 ± 0.100 0.894 ± 0.180 0.030 ± 0.015 0.193 ± 0.037 0.007 ± 0.003
Llama2-7B 3.190 ± 0.783 0.259 ± 0.184 0.769 ± 0.194 0.042 ± 0.030 0.168 ± 0.042 0.009 ± 0.006
Llama3.1-8B-it 6.003 ± 1.782 0.187 ± 0.127 1.082 ± 0.244 0.028 ± 0.019 0.235 ± 0.055 0.006 ± 0.004
Llama3.3-70B-it 4.883 ± 1.106 0.112 ± 0.093 0.942 ± 0.198 0.016 ± 0.012 0.206 ± 0.043 0.003 ± 0.003
Nemotron3.1-it 5.125 ± 1.284 0.114 ± 0.094 0.985 ± 0.207 0.016 ± 0.012 0.216 ± 0.046 0.003 ± 0.003
Phi3-Medium-128K-it 5.063 ± 1.914 0.584 ± 0.559 1.000 ± 0.319 0.087 ± 0.083 0.217 ± 0.068 0.019 ± 0.018
Phi4 5.929 ± 1.545 0.657 ± 0.472 1.306 ± 0.231 0.103 ± 0.070 0.286 ± 0.050 0.022 ± 0.015
Mistral0.3-7B 4.718 ± 1.340 0.398 ± 0.405 0.941 ± 0.240 0.059 ± 0.059 0.206 ± 0.053 0.013 ± 0.013
Qwen2.5-14B 5.184 ± 2.241 1.270 ± 1.547 1.005 ± 0.288 0.182 ± 0.221 0.223 ± 0.067 0.040 ± 0.052
DeepSeekR1L-8B 5.502 ± 1.549 0.189 ± 0.118 0.955 ± 0.204 0.028 ± 0.017 0.209 ± 0.046 0.006 ± 0.004
DeepSeekR1Q-14B 5.126 ± 2.375 1.406 ± 1.609 0.900 ± 0.269 0.198 ± 0.226 0.199 ± 0.062 0.044 ± 0.052
Table 5: Quantization error (mean ± std) for the key and value caches ($K_{i}$ and $V_{i}$) at 2-bit, 3-bit, and 4-bit quantization, evaluated on the C4 dataset.
Model K2 V2 K3 V3 K4 V4
Llama3.2-1B 4.885 ± 1.056 0.207 ± 0.166 1.074 ± 0.289 0.030 ± 0.024 0.233 ± 0.062 0.006 ± 0.005
Llama3.2-1B-it 4.524 ± 1.108 0.193 ± 0.137 0.925 ± 0.235 0.028 ± 0.020 0.201 ± 0.050 0.006 ± 0.005
Llama3.2-3B 3.885 ± 0.777 0.282 ± 0.150 0.909 ± 0.168 0.042 ± 0.023 0.194 ± 0.032 0.009 ± 0.005
Llama3.2-3B-it 4.135 ± 1.088 0.274 ± 0.137 0.912 ± 0.176 0.039 ± 0.020 0.195 ± 0.036 0.009 ± 0.004
Llama2-7B 6.337 ± 1.710 0.456 ± 0.247 1.054 ± 0.263 0.071 ± 0.038 0.213 ± 0.052 0.015 ± 0.008
Llama3.1-8B-it 6.262 ± 1.789 0.254 ± 0.185 1.128 ± 0.249 0.036 ± 0.026 0.247 ± 0.056 0.008 ± 0.005
Llama3.3-70B-it 4.391 ± 1.027 0.121 ± 0.097 0.847 ± 0.175 0.017 ± 0.013 0.186 ± 0.038 0.004 ± 0.003
Nemotron3.1-it 5.367 ± 1.332 0.127 ± 0.105 1.049 ± 0.222 0.018 ± 0.013 0.231 ± 0.049 0.004 ± 0.003
Phi3-Medium-128K-it 4.831 ± 1.759 0.788 ± 0.726 1.022 ± 0.306 0.109 ± 0.097 0.220 ± 0.064 0.023 ± 0.021
Phi4 5.715 ± 1.442 0.850 ± 0.684 1.316 ± 0.245 0.124 ± 0.093 0.291 ± 0.056 0.027 ± 0.020
Mistral0.3-7B 5.027 ± 1.332 0.543 ± 0.493 1.014 ± 0.269 0.079 ± 0.068 0.223 ± 0.060 0.017 ± 0.015
Qwen2.5-14B 4.382 ± 2.170 1.544 ± 1.872 0.846 ± 0.250 0.220 ± 0.265 0.187 ± 0.060 0.048 ± 0.060
DeepSeekR1L-8B 4.575 ± 1.122 0.204 ± 0.134 0.817 ± 0.141 0.030 ± 0.019 0.179 ± 0.033 0.006 ± 0.004
DeepSeekR1Q-14B 4.832 ± 2.354 1.651 ± 1.914 0.927 ± 0.283 0.232 ± 0.267 0.201 ± 0.061 0.051 ± 0.060
Table 6: Quantization error (mean ± std) for the key and value caches ($K_{i}$ and $V_{i}$) at 2-bit, 3-bit, and 4-bit quantization, evaluated on the GSM8K dataset.
Model K2 V2 K3 V3 K4 V4
Llama3.2-1B 5.703 ± 1.557 0.179 ± 0.136 1.213 ± 0.352 0.026 ± 0.020 0.266 ± 0.078 0.005 ± 0.004
Llama3.2-1B-it 5.002 ± 1.383 0.171 ± 0.130 1.024 ± 0.287 0.025 ± 0.020 0.223 ± 0.061 0.006 ± 0.004
Llama3.2-3B 4.840 ± 1.396 0.261 ± 0.136 1.045 ± 0.211 0.038 ± 0.021 0.226 ± 0.044 0.008 ± 0.005
Llama3.2-3B-it 3.604 ± 0.850 0.226 ± 0.129 0.790 ± 0.135 0.034 ± 0.019 0.171 ± 0.028 0.007 ± 0.004
Llama2-7B 5.081 ± 1.396 0.405 ± 0.231 0.969 ± 0.238 0.065 ± 0.037 0.205 ± 0.050 0.014 ± 0.008
Llama3.1-8B-it 6.445 ± 1.837 0.213 ± 0.161 1.184 ± 0.268 0.030 ± 0.022 0.257 ± 0.060 0.007 ± 0.005
Llama3.3-70B-it 4.967 ± 1.127 0.113 ± 0.091 0.978 ± 0.203 0.016 ± 0.012 0.214 ± 0.044 0.004 ± 0.003
Nemotron3.1-it 4.752 ± 1.124 0.113 ± 0.089 0.940 ± 0.194 0.016 ± 0.012 0.206 ± 0.042 0.004 ± 0.003
Phi3-Medium-128K-it 4.940 ± 1.834 0.605 ± 0.579 1.042 ± 0.320 0.088 ± 0.082 0.227 ± 0.069 0.019 ± 0.018
Phi4 6.610 ± 1.624 0.785 ± 0.598 1.498 ± 0.293 0.116 ± 0.082 0.330 ± 0.064 0.025 ± 0.017
Mistral0.3-7B 5.308 ± 1.367 0.461 ± 0.434 1.065 ± 0.288 0.067 ± 0.061 0.232 ± 0.061 0.015 ± 0.013
Qwen2.5-14B 4.829 ± 2.179 1.736 ± 2.659 0.979 ± 0.264 0.241 ± 0.372 0.214 ± 0.061 0.051 ± 0.077
DeepSeekR1L-8B 5.547 ± 1.517 0.193 ± 0.129 1.000 ± 0.212 0.028 ± 0.018 0.218 ± 0.049 0.006 ± 0.004
DeepSeekR1Q-14B 4.477 ± 2.176 1.424 ± 1.752 0.830 ± 0.256 0.200 ± 0.242 0.181 ± 0.058 0.044 ± 0.056
Figure 8: Mean squared error (MSE) of KV quantization for Llama 3.3-70B on C4. We use quantization bit-widths ranging from 2 to 8. The x-axis represents the quantization bit-width, while the y-axis shows the MSE on a logarithmic scale. A logarithmic scale is used to highlight differences at higher bit-widths, where MSE values decrease significantly and approach zero, particularly at 8-bit. Solid lines indicate the mean MSE across layers, while the shaded regions represent the min-max range of errors.
(a) C4 dataset
(b) MMLU dataset
(c) GSM8k dataset
Figure 9: Quantization error (MSE) of K and V caches in the Llama 3.1 8B model across 32 layers for the (a) C4, (b) MMLU, and (c) GSM8k datasets. The top row in each plot shows 2-bit quantization (k2, v2), while the bottom row shows 4-bit quantization (k4, v4). Each point represents a different attention head in the corresponding layer. The x-axis indicates the layer index, and the y-axis shows the quantization error in MSE.
(a) C4 dataset
(b) MMLU dataset
(c) GSM8k dataset
Figure 10: Quantization error (MSE) of K and V caches in the Llama 3.1 8B model across 32 layers for the (a) C4, (b) MMLU, and (c) GSM8k datasets. The top row in each plot shows 2-bit quantization (k2, v2), while the bottom row shows 4-bit quantization (k4, v4). Each point represents a different attention head in the corresponding layer. The x-axis indicates the layer index, and the y-axis shows the quantization error in MSE.
Appendix G Reproducibility and Resources
Inference was performed using the Hugging Face Transformers [53] and Accelerate [54] libraries, together with FlashAttention [55, 56], to optimize token-generation throughput. We integrated both Quanto and HQQ into the Language Model Evaluation Harness [57] to enable systematic and reproducible evaluation of model performance under varying quantization schemes. All runs were executed on GPU clusters equipped with NVIDIA A100 (80 GB) using the Quanto backend and NVIDIA H200 (141 GB) using the HQQ backend.
Appendix H Extended Evaluation Results
Extended evaluation results for KV cache quantization under the HQQ backend are presented in Tables 910 (GSM8K flexible and strict accuracies), Tables 8 (EQ-Parseable parseable-pass rates and exact-match scores), and Table 7 (CoQA F1 scores).
Table 7: Full CoQA Accuracies F1.
Variant Qwen3-0.6B Llama-3.2-1B-Instruct Llama-3.2-3B-Instruct Qwen3-4B Llama-3.1-8B-Instruct Qwen3-8B Qwen3-14B Qwen3-32B
k1v1 0.129657 0.147672 0.127875 0.222218 0.127204 0.358652 0.272695 0.231077
k1v2 0.226803 0.236351 0.222816 0.360911 0.206763 0.384316 0.329965 0.327598
k1v4 0.226441 0.236751 0.229168 0.375805 0.241965 0.385504 0.323921 0.34914
k1v6 0.214566 0.242716 0.229515 0.378638 0.220753 0.378799 0.34787 0.351566
k1v8 0.219723 0.240167 0.221966 0.364095 0.241671 0.382309 0.341797 0.3545
k2v1 0.252746 0.189839 0.576534 0.463689 0.565141 0.598736 0.73744 0.732328
k2v2 0.334262 0.525527 0.760021 0.72138 0.757424 0.766938 0.801972 0.808949
k2v4 0.333106 0.558613 0.765639 0.732337 0.762769 0.760775 0.799323 0.808594
k2v6 0.339966 0.564876 0.764105 0.730226 0.763797 0.765849 0.797442 0.80742
k2v8 0.331698 0.567622 0.763972 0.733084 0.765828 0.769881 0.79908 0.803599
k4v1 0.496626 0.575004 0.768665 0.803298 0.773274 0.815164 0.823299 0.817609
k4v2 0.666081 0.693266 0.797161 0.805846 0.785901 0.822081 0.819957 0.819761
k4v4 0.691803 0.701052 0.797872 0.806737 0.788327 0.818848 0.816583 0.824813
k4v6 0.688022 0.701648 0.797325 0.809305 0.784273 0.817312 0.818533 0.823784
k4v8 0.686547 0.700206 0.79702 0.807328 0.782431 0.817872 0.818202 0.822701
k6v1 0.649633 0.597133 0.770615 0.801156 0.765218 0.812864 0.817002 0.816527
k6v2 0.692575 0.699732 0.798369 0.80343 0.784922 0.820546 0.821854 0.825656
k6v4 0.701547 0.705158 0.79613 0.806657 0.786 0.820736 0.822193 0.824622
k6v6 0.700741 0.705549 0.795475 0.807525 0.790379 0.81943 0.821026 0.825455
k6v8 0.693932 0.702742 0.79642 0.806614 0.789155 0.81943 0.820462 0.825005
k8v1 0.650554 0.598797 0.768704 0.802301 0.767131 0.814654 0.818909 0.818461
k8v2 0.691099 0.696408 0.798332 0.802632 0.787855 0.822552 0.822365 0.824697
k8v4 0.70116 0.703779 0.794958 0.807328 0.788884 0.820734 0.822508 0.825976
k8v6 0.698706 0.70461 0.795362 0.808917 0.786832 0.818963 0.820002 0.825695
k8v8 0.696944 0.70377 0.794663 0.808457 0.785888 0.820653 0.819874 0.826971
bf16 0.707878 0.70117 0.794959 0.807871 0.785389 0.820364 0.81985 0.826088
Table 8: Full EQ-Bench Parseable.
Variant Qwen3-0.6B Llama-3.2-1B-Instruct Llama-3.2-3B-Instruct Qwen3-4B Llama-3.1-8B-Instruct Qwen3-8B Qwen3-32B
k1v1 0 0 0 0 0 0 0
k1v2 0 0 0 0 0 0 0
k1v4 0 0 0 0 0 0 0
k1v6 0 0 0 0 0 0 0
k1v8 0 0 0 0 0 0 0
k2v1 0 0 8.187 0 35.673 0 2.339
k2v2 0 16.374 80.702 27.485 85.965 65.497 67.836
k2v4 0 34.503 84.795 39.766 88.889 70.175 70.175
k2v6 0 39.181 83.041 36.842 90.059 67.836 69.591
k2v8 0 38.012 84.211 40.936 88.889 68.421 70.760
k4v1 0 47.953 58.480 77.778 83.626 98.830 77.778
k4v2 63.743 97.076 95.322 87.719 97.661 95.322 80.702
k4v4 65.497 97.076 98.830 84.795 99.415 95.322 80.702
k4v6 64.328 97.076 98.830 85.380 98.830 93.567 80.702
k4v8 62.573 97.076 98.830 84.795 98.830 93.567 80.702
k6v1 23.392 60.234 53.801 81.287 90.643 96.491 79.532
k6v2 99.415 97.076 94.737 87.719 97.661 94.737 80.702
k6v4 99.415 97.661 99.415 87.135 98.830 95.906 80.702
k6v6 99.415 97.661 98.830 86.550 98.830 96.491 80.702
k6v8 99.415 97.661 98.830 87.135 98.830 97.661 80.702
k8v1 26.901 60.234 56.725 78.947 92.983 97.076 77.193
k8v2 99.415 97.076 94.737 87.135 97.661 96.491 79.532
k8v4 99.415 97.661 99.415 87.719 98.830 95.322 80.702
k8v6 99.415 97.661 99.415 86.550 98.830 95.906 80.702
k8v8 99.415 97.661 99.415 86.550 98.830 95.906 80.702
bf16 100 97.661 99.415 87.719 98.830 94.737 79.532
Table 9: Full GSM8K Accuracies (Flexible).
Variant Qwen3-0.6B Llama-3.2-1B-Instruct Llama-3.2-3B-Instruct Qwen3-4B Llama-3.1-8B-Instruct Qwen3-8B Qwen3-14B Qwen3-32B
k1v1 0.0106141 0.0121304 0.0159212 0.00985595 0.0159212 0.0106141 0.0144049 0.0204701
k1v2 0.00833965 0.0144049 0.015163 0.00985595 0.0197119 0.0128886 0.0174375 0.0181956
k1v4 0.0090978 0.0197119 0.025019 0.0144049 0.0204701 0.0075815 0.00985595 0.0106141
k1v6 0.00985595 0.0106141 0.0272934 0.0174375 0.0197119 0.0075815 0.0045489 0.0159212
k1v8 0.00833965 0.015163 0.0280516 0.0090978 0.015163 0.0136467 0.0113723 0.0106141
k2v1 0.0030326 0.0181956 0.015163 0.00833965 0.0500379 0.0128886 0.0204701 0.0144049
k2v2 0.0030326 0.0288097 0.278241 0.0265353 0.38514 0.128886 0.327521 0.38514
k2v4 0.0030326 0.0197119 0.351024 0.0432146 0.45489 0.170584 0.412434 0.43442
k2v6 0.00379075 0.0257771 0.3442 0.040182 0.445792 0.169067 0.396513 0.419257
k2v8 0.00379075 0.0280516 0.354814 0.0348749 0.460197 0.166793 0.394238 0.448825
k4v1 0.0227445 0.0727824 0.281274 0.576952 0.517058 0.77862 0.70887 0.721001
k4v2 0.040182 0.294162 0.639121 0.808188 0.756634 0.87718 0.804397 0.648976
k4v4 0.040182 0.332828 0.652767 0.830174 0.777862 0.87718 0.823351 0.629265
k4v6 0.055345 0.33207 0.66793 0.831691 0.762699 0.88097 0.828658 0.619409
k4v8 0.0416983 0.341168 0.658074 0.826384 0.75815 0.877938 0.835481 0.631539
k6v1 0.0932525 0.0962851 0.307809 0.632297 0.521607 0.795299 0.742987 0.701289
k6v2 0.367703 0.3116 0.644428 0.830174 0.761941 0.880212 0.824867 0.59818
k6v4 0.404852 0.333586 0.653525 0.836998 0.770281 0.877938 0.827142 0.610311
k6v6 0.420015 0.336619 0.662623 0.843821 0.771797 0.884003 0.8279 0.595906
k6v8 0.41395 0.341926 0.664898 0.840788 0.771039 0.886277 0.826384 0.608036
k8v1 0.0879454 0.0955269 0.304018 0.636846 0.529947 0.797574 0.742987 0.708112
k8v2 0.389689 0.313116 0.64746 0.825625 0.74981 0.882487 0.822593 0.605762
k8v4 0.416224 0.340409 0.660349 0.832449 0.770281 0.874905 0.831691 0.620925
k8v6 0.419257 0.341926 0.667172 0.843063 0.767248 0.880212 0.839272 0.608036
k8v8 0.413192 0.333586 0.666414 0.843063 0.758908 0.880212 0.839272 0.603487
bf16 0.412434 0.328279 0.663381 0.845337 0.759666 0.87718 0.835481 0.603487
Table 10: Full GSM8K Accuracies (Strict).
Variant Qwen3-0.6B Llama-3.2-1B-Instruct Llama-3.2-3B-Instruct Qwen3-4B Llama-3.1-8B-Instruct Qwen3-8B Qwen3-14B Qwen3-32B
k1v1 0 0 0 0 0 0 0 0
k1v2 0 0 0 0 0 0 0 0
k1v4 0 0 0 0 0 0 0.00075815 0
k1v6 0 0 0.0015163 0 0 0.00075815 0 0
k1v8 0 0 0 0 0 0 0.00075815 0
k2v1 0 0.00075815 0.00530705 0 0.0280516 0 0.00075815 0.0015163
k2v2 0 0.0113723 0.275967 0.00379075 0.379833 0.0667172 0.260045 0.234268
k2v4 0 0.00985595 0.350265 0.0090978 0.443518 0.103867 0.343442 0.334344
k2v6 0 0.0121304 0.3442 0.00985595 0.438969 0.115239 0.348749 0.329795
k2v8 0 0.0136467 0.353298 0.00530705 0.448825 0.109174 0.35254 0.347233
k4v1 0.0060652 0.0614102 0.292646 0.64746 0.500379 0.785444 0.824867 0.701289
k4v2 0.0424564 0.297195 0.627748 0.824867 0.750569 0.875663 0.882487 0.726308
k4v4 0.0439727 0.331312 0.642911 0.843821 0.758908 0.875663 0.877938 0.733889
k4v6 0.055345 0.331312 0.656558 0.839272 0.747536 0.880212 0.887036 0.733131
k4v8 0.0485216 0.341926 0.648976 0.834723 0.745262 0.87718 0.887036 0.736922
k6v1 0.0833965 0.0826384 0.326005 0.698256 0.505686 0.802881 0.830174 0.702047
k6v2 0.369219 0.307809 0.639121 0.839272 0.74981 0.879454 0.884003 0.70887
k6v4 0.404852 0.333586 0.643669 0.844579 0.755876 0.876422 0.884761 0.718726
k6v6 0.421531 0.335102 0.6558 0.852919 0.751327 0.881729 0.884003 0.71721
k6v8 0.41395 0.341168 0.657316 0.84837 0.752843 0.884761 0.884003 0.721001
k8v1 0.0849128 0.0841547 0.32373 0.704321 0.512509 0.805155 0.829416 0.695224
k8v2 0.388173 0.313874 0.639879 0.829416 0.739196 0.881729 0.886277 0.71721
k8v4 0.416983 0.340409 0.652009 0.845337 0.755876 0.874905 0.88097 0.721001
k8v6 0.419257 0.339651 0.661107 0.84837 0.749052 0.878696 0.886277 0.722517
k8v8 0.415466 0.33207 0.659591 0.850644 0.743745 0.876422 0.886277 0.723275
bf16 0.413192 0.327521 0.657316 0.852161 0.741471 0.874147 0.890826 0.717968